OpenStudy (anonymous):

−3(x + 3) = −3(x + 1) − 5 ANSWER PLEASE !

OpenStudy (anonymous):

You have to distribute the -3 first

OpenStudy (anonymous):

So it would become -3x-9=-3x-3-5 Do you know what to do from there?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

combine like terms So add -3 and -5

OpenStudy (anonymous):

OMG Just Give Me The Answer

OpenStudy (amistre64):

good bye

OpenStudy (anonymous):

No...that's not how this works

OpenStudy (anonymous):

\[-3x-9 = -3x-3-5\] like terms together: \[-3x+3x = 9-8\] what does this tell you?